# DISCOVER: Data-driven Identification of Sub-activities via Clustering and Visualization for Enhanced Recognition of Human Activities in Smart Homes

### v1

@@ Abstract.
Human Activity Recognition (HAR) using ambient sensors has great potential for practical applications, particularly in elder care and independent living. However, deploying HAR systems in real-world settings remains challenging due to the high cost of labeled data, the need for pre-segmented sensor streams, and the lack of flexibility in activity granularity. To address these limitations, we introduce DISCOVER, a method designed to discover fine-grained human sub-activities from unlabeled sensor data without relying on pre-segmentation. DISCOVER combines unsupervised feature extraction and clustering with a user-friendly visualization tool to streamline the labeling process. DISCOVER enables domain experts to efficiently annotate only a minimal set of representative cluster centroids, reducing the annotation workload to a small number of samples ($0.05\%$ of our dataset). We demonstrate DISCOVER’s effectiveness through a re-annotation exercise on widely used HAR datasets, showing that it uncovers finer-grained activities and produces more nuanced annotations than traditional coarse labels. DISCOVER represents a step toward practical, deployable HAR systems that adapt to diverse real environments.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Human-centered computing Ubiquitous and mobile computingccs: Computing methodologies Machine learning approaches
@@ 1. Introduction
Human Activity Recognition (HAR) using ambient sensors integrated into the built environment is an increasingly dynamic research field that has a potential to revolutionize a number of healthcare and monitoring use cases (Alam et al., 2012; Kientz et al., 2008a; Ransing and Rajput, 2015). Practical applications of HAR include support for independent living for seniors (Sasaki et al., 2016; Augusto et al., 2024; Riboni et al., 2015) and health monitoring (Fritz et al., 2022; Civitarese, 2017). In elder care contexts, HAR can serve as a foundation for longitudinal assessments of shifts in daily habits and behaviors. Such analyses may enable the early detection of emerging physical and cognitive challenges, therefore supporting timely interventions and improved patient outcomes (Morita et al., 2023; Riboni et al., 2016).
Deploying Human Activity Recognition requires the instrumentation of target environments (e.g., residential spaces) with multimodal sensor networks, followed by the development of algorithmic approaches to transform the resultant sensor data streams into meaningful activity labels. Existing HAR approaches commonly employ fully supervised learning paradigms (Liciotti et al., 2020; Bouchabou et al., 2021b; Li et al., 2019; Demirel et al., 2018). Several large-scale, annotated datasets have been established to facilitate these approaches, including CASAS (Cook et al., 2009), Marble (Arrotta et al., 2022), and Orange4Home (Cumin et al., 2017).
The ultimate goal of HAR research is to develop techniques that are practical and deployable in a novel real-world setting. Toward this goal, we argue that successful methodologies should have the following characteristics:
@@
    Minimal dependence on labeled data: Obtaining large quantities of labeled training data for each new environment is prohibitively expensive as data collection and annotation in real homes remains both resource-intensive and logistically challenging. Sensor equipment must be installed for many months to capture diverse human behaviors, and the resulting large-scale sensor logs require painstaking and costly annotation by trained experts or residents. HAR methodologies should therefore either support generalization and transfer between environments (Thukral et al., 2024), be sample efficient, such as by utilizing active learning (Hiremath et al., 2022; Hiremath and Ploetz, 2024), or both.
@@
    Independence from prior segmentation: Numerous algorithms (Thukral et al., 2024; Hiremath et al., 2022; Demirel et al., 2018) operate under the assumption that sensor streams are pre-segmented into discrete, single-activity intervals. However, the temporal segmentation of continuous activity data presents significant challenges in delineating precise activity boundaries. Therefore, robust HAR approaches should not presuppose the availability of such segmentation.
@@
    Flexible activity granularity: There is no consensus in existing HAR literature regarding the optimal definition and scope of an ”activity,” in large part because activity relevance and scope varies substantially across application domains. While some use cases may prioritize macro-level activities (e.g., sleep, cooking), others require fine-grained activity recognition (e.g., walking from fridge to stove). Current supervised learning approaches, including those supporting cross-domain transfer (Thukral et al., 2024), are constrained by the predefined coarse activity labels within their training datasets.
In this work, we introduce DISCOVER, a self-supervised system designed to discover fine-grained human sub-activities from unlabeled sensor data without relying on pre-segmentation. DISCOVER advances the field in two complimentary ways. First, we contribute a novel HAR approach that uniquely satisfies all three of the above requirements. Second, we contribute an open source tool that facilitates a shift in how HAR annotation is conducted, enabling the dynamic discovery of finer-grained sub-activities, accommodating a wider variety of behaviors observed in real homes.
We propose the use of a minimal-effort, human-in-the-loop annotation workflow, with domain experts only needing to label a small number of representative cluster centroids, corresponding to only $0.05\%$ of our overall dataset. These labeled centroids are then used to automatically propagate annotations to other similar data points. This design substantially reduces the manual labeling burden while still preserving interpretability and control over the final labels. Beyond bypassing the label scarcity problem, this methodology also relaxes the pre-segmentation assumption, allowing continuous time-series data to be naturally segmented based on the sensor readings themselves.
We demonstrate the effectiveness of our approach through a re-annotation exercise on widely used HAR datasets (Cook et al., 2009). By applying our pipeline to an already labeled dataset, we show that our method can uncover finer-grained activities and yield more nuanced, contextually relevant annotations than the original coarse labels. Our key contributions can be summarized as follows:
@@
    Self-Supervised Pipeline for Human Sub-Activity Discovery: We design a clustering-based approach that learns directly from raw sensor streams—no large, pre-annotated training sets required—revealing the underlying sub-activities, often overlooked by conventional methods.

### v3

# DISCOVER: Identifying Patterns of Daily Living in Human Activities from Smart Home Data

@@ Abstract.
Smart homes equipped with ambient sensors offer a transformative approach to continuous health monitoring and assisted living. Traditional research in this domain primarily focuses on Human Activity Recognition (HAR), which relies on mapping sensor data to a closed set of predefined activity labels. However, the fixed granularity of these labels often constrains their practical utility, failing to capture the subtle, household-specific nuances essential, for example, for tracking individual health over time. To address this, we propose DISCOVER, a framework for discovering and annotating Patterns of Daily Living (PDL)—fine-grained, recurring sequences of sensor events that emerge directly from a resident’s unique routines. DISCOVER utilizes a self-supervised feature extraction and representation-aware clustering pipeline, supported by a custom visualization interface that enables experts to interpret and label discovered patterns with minimal effort. Our evaluation across multiple smart-home environments demonstrates that DISCOVER identifies cohesive behavioral clusters with high inter-rater agreement while achieving classification performance comparable to fully-supervised baselines using only $0.01\%$ of the labels. Beyond reducing annotation overhead, DISCOVER establishes a foundation for longitudinal analysis. By grounding behavior in a resident’s specific environment rather than rigid semantic categories, our framework facilitates the observation of within-person habitual drift. This capability positions the system as a potential tool for identifying subtle behavioral indicators associated with early-stage cognitive decline in future longitudinal studies.
@@
copyright: acmlicensedjournalyear: 2026ccs: Human-centered computing Ubiquitous and mobile computingccs: Computing methodologies Machine learning approaches
@@ 1. Introduction
What is the precise definition of a human activity in the context of smart homes? Currently, no consensus exists in the literature, primarily because the relevance and scope of an activity are inherently dependent on the specific application (Dhekane and Ploetz, 2025b). A large portion of research studying human activities in smart homes views it as a human activity recognition (HAR) problem (Ahmed et al., 2022; Demirel et al., 2018; Cook et al., 2009). HAR promises to provide a foundation for longitudinal assessments of changes in daily habits, which can facilitate the early detection of emerging physical or cognitive challenges, enabling timely interventions and improved patient outcomes (Morita et al., 2023; Riboni et al., 2016). In this setting, HAR relies on a fixed vocabulary of activity labels, such as “Cooking”, “Bathing”, and “Sleeping”. These specific activities are usually derived from Activities of Daily Living (ADL) – the basic skills necessary for individuals to independently care for themselves (Edemekong et al., 2025). However, day-to-day behavioral patterns in a given home do not neatly conform to these standardized definitions, frequently involving personal habits and unique household routines that fall outside the scope of general-purpose activity datasets. For example, the activity of “Relaxing” across various smart homes can exhibit a wide range of different physical manifestations, such as strolling around the house, listening to music while sitting in a chair, or watching TV. Furthermore, many daily activities do not correspond to any of the predefined labels, resulting in large catch-all labels that provide no useful insights into the underlying behavior of residents.
An alternative to the traditional activity-centric HAR paradigm is the identification of Patterns of Daily Living (PDL). PDLs represent meaningful, fine-grained, and recurring sequences of sensor events that capture the data-driven structure of how residents interact with their specific environments. Rather than forcing human behavior into a limited set of predefined categories, PDLs emerge from latent regularities in sensor activations across space and time. These patterns may encompass canonical ADLs such as “Cooking”, but they also capture the finer interactions that constitute them, such as “Stove use” or “Table setting.” Furthermore, PDLs reveal transitional behaviors (e.g., movement from bedroom to bathroom) and highly personalized routines (such as the specific use of a particular armchair) that lack a distinct semantic equivalent in traditional activity taxonomies. By uncovering these patterns directly from the data, smart home systems can reveal the idiosyncratic organization of daily life in unique environments and, crucially, quantify how these routines evolve or deviate over time.
This distinction between predefined labels and PDLs carries significant practical implications for both modeling and practical interpretation. Traditional activity-centric approaches prioritize alignment with a closed set of annotations and are typically evaluated using standard classification metrics such as accuracy and F1 score. However, these metrics offer little insight into the underlying structure of those activities or the unique, user-specific habits manifested within a given environment. In contrast, pattern-centric approaches emphasize uncovering latent structures within continuous, unsegmented sensor data, even when those structures subdivide or intersect with traditional activity labels. While PDLs can be mapped back to canonical ADLs for benchmarking, their primary value lies in providing a flexible, high-resolution representation of daily life, which can facilitate personalized monitoring and longitudinal sensitivity analysis.
In this work, we build upon this pattern-centric viewpoint by proposing DISCOVER, a framework designed to discover PDLs from smart home data without requiring prior annotations or pre-segmentation. DISCOVER identifies these patterns through self-supervised pre-training with a masked modeling objective, generating semantically rich sensor data embeddings. These representations are further refined using a clustering objective to produce distinct clusters that correspond to unique Patterns of Daily Living. To facilitate the interpretation of these clusters, we provide a user-friendly visualization tool that replays sensor activation sequences directly onto a smart home layout, enabling a seamless labeling process accessible to non-experts. By annotating as little as $0.01\%$ of the data samples, DISCOVER propagates labels across all discovered patterns, streamlining the entire annotation process. Crucially, by removing the need for manual pre-segmentation, our approach allows continuous sensor streams to organize naturally according to their inherent behavioral structure while preserving both interpretability and user control.
The main contributions of our work are as follows:
@@
    Data-Grounded Pattern Discovery: We develop a self-supervised feature extraction and clustering pipeline that learns directly from raw, continuous sensor streams without the need to rely on large annotated training datasets to uncover data-driven patterns of daily living. Rather than fitting the data to a closed set of predefined activities, the system discovers environment-specific behavioral structures grounded in the sensor data itself.
@@
    Eliminating Pre-Segmentation Constraints: By handling continuous, unsegmented time-series, our framework is applicable to realistic smart home scenarios and can be easily extended to real-world deployments.
@@
    Efficient Human-in-the-Loop Annotation: Leveraging an active-learning-like strategy, we focus manual annotation effort on a sparse set of representative cluster centroids. By propagating labels from a small set of representative cluster centroids, DISCOVER significantly reduces the manual annotation burden. In our experiments, we achieved high classification performance by annotating only a handful of examples per cluster, representing less than $0.01\%$ of the total dataset.
@@
    Open-Source PDL Visualization Tool: We introduce an interactive, open-source visualization tool that enables researchers and practitioners to explore discovered Patterns of Daily Living, inspect their spatial and temporal structure, and relate them back to sensor layouts and canonical activities. The tool supports qualitative interpretation, providing a practical bridge between unsupervised pattern discovery and validation.
To validate our approach, we apply DISCOVER to a widely used smart-home dataset spanning multiple smart home environments. Our analysis demonstrates that DISCOVER yields coherent behavioral clusters that align with canonical activities while uncovering fine-grained structures beyond current annotation schemas. We demonstrate the practical value of this framework by showing that it achieves classification performance comparable to fully-supervised baselines with orders of magnitude fewer labels, while simultaneously enabling the detection of subtle, longitudinal habitual shifts. Crucially, our findings indicate that these discovered patterns provide a higher-resolution representation of daily life than traditional HAR labels. By anchoring behavior in a resident’s specific environment, DISCOVER bridges the gap between raw sensor data and the nuanced longitudinal insights required for future research into early-stage cognitive decline.
2. Background & Related Work
DISCOVER introduces a novel perspective for analyzing smart home data through the discovery of Patterns of Daily Living (PDL). This approach looks beyond traditional ground truth annotations, which typically relate to only a limited subset of activity granularities and often obscure the underlying behavioral motifs. By shifting the focus from predefined labels to data-grounded discovery, we establish a more flexible framework for understanding longitudinal habits.
We structure our review of related work as follows. First, Section 2.1 defines the hierarchical spectrum of activity granularity. We then survey benchmark smart home datasets through the lens of sensing density and its impact on the activity taxonomy in Section 2.2. In Section 2.3, we evaluate existing visualization and annotation methods, highlighting the reliability gaps in standard labeling protocols. Section 2.4 provides an overview of the traditional smart home HAR pipeline and its current limitations. Finally, Section 2.5 reviews the state-of-the-art in self-supervised feature extraction and clustering, establishing the technical foundation for the DISCOVER framework.
2.1. Granularity Levels of Human Activities
Human activity is not a monolithic concept but rather a hierarchical spectrum ranging from primitive physical movements to complex, longitudinal behaviors and habits (Bobick, 1997; Nagel, 1988; Vrigkas et al., 2015). At the most granular level, behaviors are composed of movements or gestures, comprised of primitive, atomic, and recurring motion units. For example, the sequential movements of lifting a leg, taking a stride, and pushing the ground backward constitute the more complex event of taking a step. When performed within a specific context, these event sequences generate activities (e.g., “Walking”). Finally, a high-level aggregation of activities denotes a behavior or a habit, a more abstract construct that typically requires longitudinal analysis to identify.
This taxonomy has been extensively studied in computer vision, where capturing micro-level gestures and environmental context from video is relatively straightforward (Bukht et al., 2025; Bobick, 1997; Nagel, 1988; Arshad et al., 2022). However, achieving a similar resolution using sparse modalities, such as ambient smart-home sensors, remains a significant challenge. While existing work has explored identifying activity patterns in smart environments (Boovaraghavan et al., 2023; Bourobou and Yoo, 2015; Alia and Lago, 2025; Yassine et al., 2017; Ni et al., 2015), the interpretation of these patterns varies wildly, ranging from high-level behavioral trends to mid-level actions.
In contrast, DISCOVER assumes having access to only ambient sensor data and is designed to operate specifically at the fine-grained pattern level, identifying the recurring motifs that bridge the gap between raw sensor triggers and semantic activities, without using any ground truth activity annotations. We distinguish DISCOVER from traditional approaches by three core guarantees. First, the discovery is grounded in the actual data: patterns emerge directly from latent regularities in the sensor stream rather than conforming to predefined, top-down categories. Second, each discovered cluster represents a single, semantically consistent routine (e.g., a specific path between the dining table and the kitchen). Third, these patterns are identified without any prior labels, bypassing the primary bottleneck of supervised HAR.
2.2. Sensing Density and Activity Taxonomy
The granularity of activity labels in smart home research is fundamentally constrained by the density and modality of the underlying sensor deployment. As established in Section 2.1, human behavior exists on a hierarchical spectrum of movements, actions, activities, and behaviors. Existing smart home datasets typically occupy specific tiers of this hierarchy based on the instrumentation of their respective deployments.
Most widely-used benchmarks, such as CASAS (Cook et al., 2012), Van Kasteren (Van Kasteren et al., 2010), and ARAS (Alemdar et al., 2013), represent sparse ambient sensing environments. These deployments typically utilize motion and contact sensors limited to one to two per room or primary appliance (De-La-Hoz-Franco et al., 2018). Consequently, these datasets are primarily annotated at the activity level, capturing canonical ADLs like Cooking, Sleeping, or Personal Hygiene. While these labels are semantically clear, they often collapse heterogeneous behaviors into single, non-descriptive categories. For example, a Relax label may encompass reading in an armchair, watching television, or napping. The sparse nature of these deployments inherently limits the tracking of the finer-grained actions required to distinguish these sub-routines.
Conversely, datasets like Placelab (Intille et al., 2005) and the MIT activity recognition datasets (Tapia et al., 2004) utilize significantly denser instrumentation, often including hundreds of sensors to monitor fine-grained interactions. These datasets provide annotations at the action or movement granularity. While these environments offer a rich view of human-object interaction, such extensive instrumentation is difficult to scale to real-world residential settings due to the prohibitive complexity and maintenance of the hardware (Boovaraghavan et al., 2023).
Rather than being restricted by a specific sensing modality, DISCOVER is designed to be configuration-adaptive. It does not rely on pre-defined activity mappings that may not be supported by the available instrumentation. Instead, DISCOVER decomposes available data into meaningful Patterns of Daily Living (PDLs). DISCOVER is capable of identifying localized sub-routines, such as specific trajectories through a kitchen or stationary seating preferences, that standard HAR pipelines typically discard as non-descriptive noise. This approach allows DISCOVER to extract high-resolution behavioral insights from existing sparse datasets without requiring the dense instrumentation of lab-based deployments.
2.3. Smart Home Visualization and Annotation Methods
The hierarchical temporal patterns inherent in smart home data require robust visualization and annotation protocols. While specific actions, such as taking medicine, can be resolved via simple pillbox sensors, broader activities like Cooking often comprise a stochastic sequence of sub-activities (e.g., pantry access, food preparation, and stove usage) captured only through ambient motion triggers. Furthermore, the temporal disparity between long-duration events, such as sleeping, and near-instantaneous transitions, such as leaving home, complicates the task of consistent data interpretation.
Visualization in the current literature is frequently relegated to “displaying insights” rather than facilitating raw data exploration (Castelli et al., 2017; Le et al., 2014). Existing tools for the CASAS datasets primarily utilize spatial video overlays (Karpekov et al., 2025) and heatmaps (Chen and Dawadi, 2011) to provide context for temporal sensor activation patterns. However, the annotation protocols associated with these visualizations remain a significant bottleneck. Standard methods often rely on scheduled routines (Arrotta et al., 2022; Cumin et al., 2017), which lack the naturalistic, “in-the-wild” component, or retrospective self-reporting (Alemdar et al., 2013), which is notoriously prone to recall bias and high participant (Tonkin et al., 2018). Alternatively, manual post-hoc labeling of sensor logs, as performed in CASAS, is labor-intensive and susceptible to systematic human error and inter-rater (Demrozi et al., 2023).
These annotation inaccuracies can fundamentally undermine the reliability of HAR benchmarks and subsequent model evaluation. DISCOVER addresses these deficiencies by automating the identification of Patterns of Daily Living (PDLs), requiring expert validation for only a minimal subset of the data. By integrating an intuitive spatial visualization via our custom-made tool (Karpekov et al., 2025), DISCOVER allows non-experts to accurately interpret and label discovered patterns, thereby establishing a scalable and more reliable ground truth than traditional manual protocols.
2.4. Activity Recognition in Smart Homes
Similar to the case with body-worn IMU sensors (Bulling et al., 2014), human activity recognition (HAR) in smart homes is based on a multi-stage processing pipeline consisting of data acquisition, pre-processing, segmentation, feature extraction, and classification. HAR in smart environments has emerged as a critical area of research, enabling the identification of Activities of Daily Living (ADLs) through ambient sensors (Bouchabou et al., 2021a; van Kasteren et al., 2010; Rafferty et al., 2017; Kientz et al., 2008). ADLs are particularly valuable for routine detection, which feeds numerous downstream applications, including health monitoring and assistive systems for independent living (Chatting, 2023; Csurka, 2017; Morita et al., 2023). Conventional approaches to HAR rely on supervised learning methods, where large amounts of labeled data are needed. These methods span a range of classic machine-learning techniques, including Random Forests (Sedky et al., 2018), Hidden Markov Models (HMMs) and Support Vector Machines (SVMs) (Cook, 2010). A majority of these approaches rely on features that are hand-crafted by domain experts, and universally assume pre-segmented data, which limits their applicability in real-world scenarios.
Advances in deep learning have expanded the capabilities of HAR systems. Deep Neural Networks (DNNs) (Alaghbari et al., 2022), Long Short-Term Memory networks ((LSTM) (Liciotti et al., 2020), and hybrid architectures with Embeddings from Language Models (ELMo) combined with BiLSTM (Bouchabou et al., 2021c) have demonstrated promising performance in activity classification tasks. Additionally, graph neural networks (GNNs) have been employed to capture spatial and temporal relationships in sensor data (Li et al., 2019; Zhou et al., 2020; Plötz and others, 2023).
Transfer learning is another promising subdomain, where a model is trained on a fully labeled source home and then used to annotate a new target dataset (Dhekane and Ploetz, 2025b). Thukral et al. (Thukral et al., 2025) achieve this by using text representations of sensor activations as an intermediate modality that allows for a cross-house transfer.
Despite these advancements, the reliance on pre-segmented data remains a significant limitation. Some studies have explored segmentation techniques, such as change-point detection algorithms, which identify abrupt changes in sensor data to infer activity boundaries (Aminikhanghahi et al., 2018; Sprint et al., 2020; Jose and Malekian, 2017; Aminikhanghahi and Cook, 2017). However, these methods often rely on heuristics and are not well-suited for capturing fine-grained activities. Moreover, these HAR approaches depend on massive labeled datasets and fixed, pre-defined labels that cannot always capture the nuanced behaviors exhibited in real-world environments. To address this challenge, our work introduces a self-supervised framework to process large unlabeled datasets, paired with clustering and active learning steps that allow to discover fine-grained patterns of daily living with minimal human annotation efforts.
2.5. Clustering Pipelines
Several recent works have explored behavior inference from unlabeled data through automated class discovery. Wu et al. (Wu et al., 2020) introduced a pipeline for clustering audio embeddings followed by one-shot labeling, while Patidar et al. proposed VAX (Patidar et al., 2023), which leverages pre-trained video and audio models to bootstrap activity recognition. Similarly, Gao et al. (Gao et al., 2024) utilize Large Language Models (LLMs) to iteratively cluster and relabel activities within pre-segmented datasets. While these systems demonstrate the utility of user-in-the-loop refinement, their reliance on high-bandwidth modalities like audio and video creates a significant gap in transferability to ambient smart-home environments. Unlike acoustic or vision-based models, which benefit from rich, pre-trained feature extractors (e.g., VGGish or VideoMAE), ambient sensor streams consist of sparse, low-dimensional triggers that lack inherent semantic descriptors. Furthermore, these approaches also require pre-segmentation and assume access to rich multi-sensor pre-training signals.
In the smart home domain, unsupervised clustering methods have been employed in HAR to group unlabeled sensor data into meaningful activity clusters (Ariza Colpas et al., 2020). Classic approaches, such as Lloyd’s clustering algorithm combined with k-Nearest Neighbors (Fahad et al., 2014) and DBSCAN (Hoque and Stankovic, 2012), have been used to identify patterns in sensor data. More recently, the SCAN framework (Gansbeke et al., 2020) has been adapted for HAR, enabling the clustering of unlabeled movement data from wearable sensors (Ahmed et al., 2022). These methods leverage large-scale sensor data to enhance activity recognition, addressing the challenges associated with obtaining annotated datasets. However, existing “cluster then classify” approaches still rely on downstream labels to match and often lack the ability to discover fine-grained, data grounded activities.
Recent work by Bock et al. (Bock et al., 2024) introduced a clustering-based annotation pipeline that reduces the need for human labeling by leveraging vision-based models for weak annotation in HAR datasets, demonstrating that clustering methods can match fully supervised classifiers while minimizing manual effort. Inspired by successes in the wearable domain, we propose a novel approach tailored for ambient sensor data in smart homes that addresses the challenges of heterogeneous sensor streams while discovering fine-grained patterns of daily living rather than merely mapping back to the original labels.
3. DISCOVER: A Self-Supervised Approach of Identifying Patterns of Daily Living in Human Activities from Smart Home Data
Figure 1. Overview of DISCOVER, a self-supervised framework for the data-grounded discovery of Patterns of Daily Living (PDL), consisting of two main stages: clustering and labeling. After (0) slicing the raw data into continuous sliding windows of sensor activations without assuming any pre-segmentation, we procced to the (1) Representation Learning phase, where a Transformer-based model is trained via masked language modeling to map windows into an embedding space, followed by a (2) clustering phase, where we use these embeddings to identify similar activity windows and fine-tune a clustering model using SCAN loss, which results in assigning all data points to $k$ clusters. We then (3) sample a handful of windows closest to each cluster centroid, replay them on 2D house layouts using our custom built visualization tool, and send these samples to a group of experts for annotation. With minimal labeling effort, we obtain the annotations of the discovered PDLs for each cluster centroid, and (4) propagate them to the rest of the data points in each cluster. These annotations are then (5) applied to the original dataset. The resulting high-resolution PDL stream can be utilized for specialized tasks, such as longitudinal monitoring of behavioral drift or training data-efficient recognition models.
To address the limitations of the label-centric paradigm in smart home research, we introduce DISCOVER, a framework for the data-grounded discovery of Patterns of Daily Living (PDL). Unlike traditional approaches that attempt to map sensor streams to a closed set of activity categories, DISCOVER identifies recurring, fine-grained routines directly from unlabeled data without requiring pre-segmentation. By combining self-supervised representation learning with an interactive refinement loop, the system extracts the idiosyncratic behavioral motifs unique to a resident’s specific environment.
We define the objective as the unsupervised discovery of latent behavioral structures in ambient smart home environments. Given a continuous time-series dataset of sensor events, we partition the stream into a collection of $n$ sliding windows $\mathcal{W}=\{W_{1},W_{2},\dots,W_{n}\}$ of length $l$ and stride $t$. Our goal is to learn a mapping $f:W_{i}\mapsto c_{k}$ where $c_{k}$ represents a specific Pattern of Daily Living (PDL) cluster discovered from the underlying distribution of the data. In this formulation, $c_{k}$ does not denote a predefined semantic category (e.g., “Cooking”); rather, it represents a data-driven routine (e.g., a specific sequence of movements between a dining area and a kitchen).
Fig. 1 gives an overview of DISCOVER, which consists of two main stages: clustering and labeling. The approach starts by pre-processing sensor data into sliding windows in step (0). We then pursue the discovery through the following five steps: